OpenStudy (anonymous):

(how will you find the equation of the given slope that contains the given point?) q(5,2),slope -2

OpenStudy (amistre64):

id simply use the "point" slope" form of the equation

OpenStudy (turingtest):

y-2=-2(x-5)

OpenStudy (amistre64):

given a slope: m and a point (Px,Py) y-Py=m(x-Px)

OpenStudy (turingtest):

sorry amistre, don't mean to step on your toes

OpenStudy (amistre64):

not a problem, its a free for all answering system ;)

OpenStudy (turingtest):

why do I give you medals, aren't you level 100?

OpenStudy (amistre64):

im past level 100 but the thing stops there

OpenStudy (turingtest):

ahhh.... ic

OpenStudy (amistre64):

53470 out of -565240 XP to go ..... the next level so far is an impossibility